Suppose that we have a fuse box containing 20 fuses, of which 5 are defective. If 2 fuses are
selected at random and removed from the box in succession without replacing the first, what is
the probability that both fuses are defective?
1
Expert's answer
2020-12-04T07:46:25-0500

The probability that both fuses are defective is

"\frac{5}{20} \times \frac{4}{19} = \frac{1}{19} = 0.0526"

Answer: 0.0526